my sister is three years older than i am .
the sum of our age is 29.
how old am i ?

Answer by stanbon(75887)

my sister is three years older than i am .
the sum of our age is 29.
how old am i ?
------
Equation:
s = i + 3
s + i = 29
-----
Substitute and solve for "i":
i + 3 + i = 29
2i = 26
i = 13 yrs
--------------
Cheers,
Stan H.